Skip to end of metadata
Go to start of metadata

EST Work Group Agenda / Minutes


Facilitator:  David Burgess

Note Taker:  David Burgess


David BurgessLabCorp
Lorraine ConstableConstable Consulting
Dave HamillHL7 International
David Johnson (DJ)HL7 International
Wayne KubickHL7 International
Laura MitterHL7 International
Elizabeth Newton

Kaiser Permanente

Brian PechKaiser Permanente
Joshua Procious

HL7 International

Andrew StatlerCerner
Michael van der ZelHL7 Netherlands
Lynn LaaksoHL7 International
Jeff BrownCigna

Agenda Topics


  • Call to Order

  • Approve Meeting Minutes

  • GForge
Lynn presented new versions of GForge and requested people start testing the new version for functionality.   Goal is to move FHIR off of GForge and into Jira prior to changing GForge versions.  On a related note, there are some GForge updates to the current version that will be installed tonight.  (Should not have any downtime involved.)

  • JIRA
Josh updated all of the JIRA and Confluence add-ons and synced Dev and Prod.

  • Confluence
Josh informed us that DJ has created 6 Ballot Process Tutorials.  They can be found at Jira Ballot Process.  At present they are draft versions.

David Johnson
  • GitHub
Nothing new to discuss

  • Website
Laura is working on stand up pages.  Discussed ability to publish RSS feeds and the ability to subscribe for JIRA and Confluence Update Notifications.

  • EA Cloud
No updates.  Asked for follow-up to see if this topic needs to remain on the agenda going forward.

  • New PSS Process Updates
Last PSS meeting changes where made involving syntax.  Currently awaiting TSC approval of the changes.  Also an approval process for PSS updates has been completed.

David Johnson
  • Amazon Hosting Topics and are moving to the new Amazon servers Friday Nov 22 after 3:00pm.  This move is a high priority due to support issues on the current servers.  Down time is expected to be around 3 to 4 hours.  Concerns include hashed passwords in the membership database, and ISS Directory Builder Issues.  This Friday was chosen do to the expected low load the following week, in the case linger issues need continued investigation.  Will try hard to avoid any rollback to existing server based upon risks associated with staying on current server.

  • Project 1370: Tooling Catalog
Current count of tools is 117, down from 133.  Process to filter through catalog continues.

  • Liaisons Update
No updates at this time.

  • Publishing

David Burgess to reach out to Brian Pech to move merger process forward.

  • Work Group Health
No updates at this time.

  • Adjourn

Supporting Documents

Outline Referenceupporting Document

Action items